Assume that females have pulse rates that are normally distributed with a mean of

mu equals 72.0μ=72.0

beats per minute and a standard deviation of

sigma equals 12.5σ=12.5

beats per minute. Complete parts​ (a) through​ (c) below.

a. If 1 adult female is randomly​ selected, find the probability that her pulse rate is less than 78 beats per minute.

b. If 25 adult females are randomly​ selected, find the probability that they have pulse rates with a mean less than 78

beats per minute.
Answer #1

Please don't hesitate to give a "thumbs up" in case you're satisfied with the answer

a. P(X<78) = P(Z< (78-72)/12.5) = P(Z< 0.48) = .6844

So, p(that her pulse rate is lower than 78bpm) = .6844

b. P(Xbar < 78) =  P(Z< (78-72)/(12.5/sqrt(25)) = P(Z< 2.4) = .9918